slove:
the answer should be 41.3 e^ j15.1

208-255·1e^(-j × 0.4013) j×2.5

j is the imaginary number

the equation is in exponential form and so must the answer.
Answer #1

J (0-1013 e 208-a55-r xa-s +30 0 9205 0-3906 Jx 2-5 Jx2.S -a6-3195] +3 99-6446 Uxa.S 2 2一2-5 = 326-3195 99-6446 2-5 2.5 he 234.32033) 一1/ (29-2573) + (10.7978) へ 588-65.08 569 = V 1703-7998 39-8578 = tan (o-269151